OpenStudy (joftheworld):

@timo86m 10x - 7y = -1 -2x -7y = -25

OpenStudy (anonymous):

thid they ask what method to use?

OpenStudy (anonymous):

did*

OpenStudy (joftheworld):

huh ?

OpenStudy (anonymous):

i assume you can use any method then since they haven't explained them. THere are substitution method, and elimination and i think combination.

OpenStudy (joftheworld):

I think for finding (x,y) you subtract ??

OpenStudy (anonymous):

10x - 7y = -1 -2x -7y = -25 what i'd do is solve for y in 10x - 7y = -1 then you'd have it in a form of y=expression then plug in that expression in the y of the bottom equation 2x-7(expression)=-25 from there you can solve for x. once you solve for x you can use that number to solve for y by pluggin in the found value of x into either bottom or top equation.

OpenStudy (anonymous):

If you didn't understand that then lets do it one at time. Solve for y in 10x - 7y = -1

OpenStudy (joftheworld):

one way I know of doing this is subtract the equations by eachother so, 10x - 7y = -1 - -2x -7y = -25

OpenStudy (anonymous):

that works also and you would actually add it to cancel the -7y

OpenStudy (anonymous):

so 10x - 7y = -1 +( -2x -7y = -25)

OpenStudy (joftheworld):

hmm.. I don't know D:

OpenStudy (anonymous):

10x - 7y = -1 +( -2x -7y = -25) first distribute the + 10x - 7y = -1 2x +7y = +25 <- now distributed now we can add just add the columns 10x - 7y = -1 2x +7y = +25 ---------------------- 12x =24 so now that u added +( -2x -7y = -25) into row 1 you got 12x =24 -2x -7y = -25 you have successfully elimanted y. 12x =24 -2x -7y = -25 from here it is easy to find x.. x=24/12= 2 so bottom row is -2(2)-7y=-25 just solve for y and you are done.

OpenStudy (joftheworld):

thanks can you help me with this one last problem ?

OpenStudy (anonymous):

sure start a new one. Close this one.